OpenStudy (jojo12g):

What is the value of the expression shown when x = –4? 6 + 3(x – 5) – (x + 3) A. –9 B. 4 C. 2 D. –20

OpenStudy (aveline):

Let's substitute x for -4 in the problems: 6+3(-4-5)-(-4+3) ...and distribute 6-12-15+4-3 =-20

OpenStudy (seratul):

Plug in -4 for x. 6 + 3(-4 - 5) - (-4 +3) = 6 + 3(-9) -1(-1)

OpenStudy (jojo12g):

thanks guys so much